OpenStudy (rowanmartinfan):

@abb0t Function: g(x) = 2x2 - 8 For x ≥ 0, the inverse function is f(x) = For x ≤ 0, the inverse function is d(x) =

OpenStudy (abb0t):

@dan815

OpenStudy (dan815):

lol he is so lazy!

OpenStudy (rowanmartinfan):

ikr can u help

OpenStudy (dan815):

lets call g(x) = y=2x^2-8 The inverse function tells us the x we put in given our y \[+or- \sqrt{(y-8)/2}=x\] now its custom to make this a function of x, as that is the variable you are passing and seeing what the unknown must have been... kinda weird to think about .. but it will make sense,,,

OpenStudy (dan815):

okay now as you can see, in the sqrt there must be a positive number so... the only time x<0 is when \[- \sqrt{(x-8)/2}=y\] and for x>0 \[\sqrt{(x-8)/2}=y\]

OpenStudy (rowanmartinfan):

yea i see

OpenStudy (dan815):

for now keep these steps in mind till you figure out everything

OpenStudy (rowanmartinfan):

oh well ill need to figure it out

OpenStudy (dan815):

if u have y=f(x)... first rearrange as x=f(y) and then switch the variable and write it as y=f(x) ------------------ for example if you have function/ y=(x+8)/2 rearrange as x=2y-8, this is a x=f(y) from now all we do is where u see why you change to x and x to change to y so y=2x-8<--- this is the inverse functioon

OpenStudy (dan815):

where you see y* you change to x and x you change to y

OpenStudy (dan815):

wait lol i feel like there is something wrong http://www.wolframalpha.com/input/?i=inverse+function+of+y%3D2x%5E2+-+8

OpenStudy (dan815):

according to this.. the inverse function is the same, for x>0 and x<0